Consider a disk with 300 tracks, and the following sequence of requested tracks 80, 150, 230, 50, 220, 50, 190, 260, 210, 270. Assume that the disk arm is located at position 60 and is moving in the up position (i.e. 1 to 300). Calculate the number of tracks crossed if using the scan, c-scan and look-scan disk scheduling algorithm.
Answer #1

Given:

disk arm is located at position 60 and is moving in the up position and

sequence of requested tracks 80, 150, 230, 50, 220, 50, 190, 260, 210, 270.

1. Scan algorithm:

Since disk arm is moving in up direction, it will process the requests having track number greater than 60- so requests for tracks 80, 150, 190, 210, 220, 260, 270 will be serviced and the disk arm will reach to right end which is track 300. So total track crossed will be 300-60 = 240 tracks. Again to process the request 50 it will move in opposite direction crossing 300-50 = 250 tracks.

Total track Crossed = 240 + 250 = 490 tracks.

2. c-Scan algorithm:

Since disk arm is moving in up direction, it will process the requests having track number greater than 60- so requests for tracks 80, 150, 190, 210, 220, 260, 270 will be serviced and the disk arm will reach to right end which is track 300. So total track crossed will be 300-60 = 240 tracks. Now instead of changing the direction of disk arm, now disk arm will be positioned to track 1 and movement will go on up direction. Now to process the request 50 it only needs to cross only 50-1=49 tracks.

So,

Total track Crossed = 240 + 49 = 289 tracks.

3. Look-Scan:

Since disk arm is moving in up direction, it will process the requests having track number greater than 60- so requests for tracks 80, 150, 190, 210, 220, 260, 270 will be serviced and the disk arm will change the direction because there is no higher track request. So total track crossed will be 270-60 = 210 tracks. Now to process requested track 50 disk arm needs to cross 270-50 = 220 track.

So the total track crossed = 210+220 = 430 tracks.